What is the future value of a $930 annuity payment over five years if interest rates are 9 percent?
1
Expert's answer
2018-07-03T05:36:08-0400
FV = P((1 + r)^n - 1)/r = 930((1 + 0.09)^5 - 1)/0.09 = $5565.78.